OpenStudy (anonymous):

The length of the diagonal of a square is 8 inches. What is the area of the square?

OpenStudy (vishweshshrimali5):

diagonal of a square = \[\sqrt{2} x\] where x = side of the square Now area os a square = \[x^2 = 2x^2/2 = (\sqrt{2} x)^2/2\] So square the diagonal and divide it by 2 to get your answer